OpenStudy (anonymous):

Solve 3x2 + 6x = 8. 1 plus or minus 2 square root 33 the quantity of negative 3 plus or minus I square root 15 all over 3 the quantity of negative 3 plus or minus square root 33 all over 3 1 plus or minus 2 i square root 15

OpenStudy (sepeario):

Okay are these 5 different questions? It's really unclear.

OpenStudy (sepeario):

For the first bit to solve 3x^2 + 6x = 8, have you heard of completing the square?

OpenStudy (anonymous):

no the bottom ones are the answer choices

OpenStudy (sepeario):

Oh right. So have you heard of completing the square?

OpenStudy (anonymous):

no :/

OpenStudy (sepeario):

It's a way to solve the trinomial like you have above. I'll show you the steps. \[3x^2+6x=8\] The first step is to move the constant (8) on the other side but that is already done. What you have to do now is to take half the coefficient of x and add it to both sides. Can you do that?

OpenStudy (anonymous):

idk how.... man i must sound dumb as hell XD

OpenStudy (anonymous):

@Sepeario

OpenStudy (sepeario):

Nah this is actually pretty hard.. but very useful when you learn it. In the term 6x, what is the coefficient of x (the number before x essentially).

OpenStudy (anonymous):

3?

OpenStudy (campbell_st):

the easiest thing to do is rewrite the equation as \[3x^2 + 6x - 8 = 0\] then use the general quadratic formula for a quadratic \[ax^2 + bx - c = 0\] the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] your question has a = 3, b - 6 and c = -8 substitute the values and calculate the answers

OpenStudy (campbell_st):

oops b = 6 a typo in the response above.

OpenStudy (sepeario):

or you could do that.. I like completing the square. if you learn it it's much quicker than substituting values into the formula.

OpenStudy (anonymous):

lol you still at this?

OpenStudy (anonymous):

must be like a homework marathon

OpenStudy (anonymous):

lol 51 questions XD @satellite73

OpenStudy (anonymous):

i hope this is like 47 at least

OpenStudy (sepeario):

what topic is this on?

OpenStudy (anonymous):

do what @campbell_st said use \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] with \[a=3,b=6,c=-8\] that is all compute carefully

OpenStudy (sepeario):

probably quadratics and polynomials and stuff like that

OpenStudy (anonymous):

i am i got \[-6\pm \sqrt{132}\6\]

OpenStudy (anonymous):

want me to check it?

OpenStudy (anonymous):

idk i just want the answer XD

OpenStudy (anonymous):

im tired of algebra. its been 3 hours and im still on the same freaking thing!

OpenStudy (anonymous):

that is the answer, before you "simplify it" it is probably not one of your answer choices

OpenStudy (anonymous):

is one of them \[\frac{-3\pm\sqrt{33}}{3}\]?

OpenStudy (anonymous):

yes XD

OpenStudy (anonymous):

then pick that one

OpenStudy (anonymous):

how many more?

OpenStudy (anonymous):

What is the simplified form of 3 over 2x minus 5 + 21 over 8 x squared minus 14x minus 15 ? 6 times the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

those are just answer choices XD

OpenStudy (anonymous):

and i think that it is 6 times the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

\[\frac{3}{2x-5}+\frac{21}{8x^2-14x-15}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i can't read your answer really i know the denominator is correct \[(2x-5)(4x+3)\]

OpenStudy (anonymous):

6(x+1)/(2x-5)(4x+3)

OpenStudy (anonymous):

What is the simplified form of the quantity 15 x y squared over x squared minus 5x plus 6 all over the quantity 5 x squared y over 2x squared minus 7x plus 3 ? 3y times the quantity 2x plus 1 over x times the quantity x plus 2 x times the quantity x plus 2 over 3y times the quantity 2x plus 1 x times the quantity x minus 2 over 3y times the quantity 2x minus 1 3y times the quantity 2x minus 1 over x times the quantity x minus 2

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

hard to read again let me try

OpenStudy (anonymous):

XDDD

OpenStudy (anonymous):

\[\frac{15xy^2}{x^2-5x+6}\]?

OpenStudy (anonymous):

i need a screenshot maybe

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

hit "print screen" on your keyboard, then attach it

OpenStudy (anonymous):

it's not working...

OpenStudy (anonymous):

ok give me a second was any part of what i wrote correct?

OpenStudy (anonymous):

i bet it was a compound fraction which amounts to \[\frac{15xy^2}{x^2-5x+6}\times \frac{2x^2-7x+3}{5x^2y}\]

OpenStudy (anonymous):